# Mastering Multi-Drone Volleyball through Hierarchical Co-Self-Play Reinforcement Learning

### v3

@@ Abstract
In this paper, we tackle the problem of learning to play 3v3 multi-drone volleyball, a new embodied competitive task that requires both high-level strategic coordination and low-level agile control. The task is turn-based, multi-agent, and physically grounded, posing significant challenges due to its long-horizon dependencies, tight inter-agent coupling, and the underactuated dynamics of quadrotors. To address this, we propose Hierarchical Co-Self-Play (HCSP), a hierarchical reinforcement learning framework that separates centralized high-level strategic decision-making from decentralized low-level motion control. We design a three-stage population-based training pipeline to enable both strategy and skill to emerge from scratch without expert demonstrations: (I) training diverse low-level skills, (II) learning high-level strategy via self-play with fixed low-level controllers, and (III) joint fine-tuning through co-self-play. Experiments show that HCSP achieves superior performance, outperforming non-hierarchical self-play and rule-based hierarchical baselines with an average 82.9% win rate and a 71.5% win rate against the two-stage variant. Moreover, co-self-play leads to emergent team behaviors such as role switching and coordinated formations, demonstrating the effectiveness of our hierarchical design and training scheme. The project page is at https://sites.google.com/view/hi-co-self-play.
@@
Keywords: Multi-Agent Reinforcement Learning, Self-Play, Robot Learning
@@ 1 Introduction
To address these challenges, we propose Hierarchical Co-Self-Play (HCSP), a hierarchical reinforcement learning (HRL) framework that separates strategic reasoning from motor control. The policy is decomposed into a high-level strategy responsible for team tactics, and multiple low-level skills responsible for drone-specific motion control. The high-level strategy is a centralized multi-layer perceptron (MLP) with three output heads, each producing commands for one drone, enabling synchronized decision-making across the team. The low-level policies are independent across drones and operate at a higher frequency to ensure responsive, fine-grained control. To reflect the turn-based structure of the game, the high-level strategy is only activated upon discrete game events–when the ball is struck or crosses the net–while low-level control runs continuously between these moments.
@@ 3.1 Stage I: Low-Level Skill Acquisition
[…]
@@
[…]
@@ 3.2 Stage II: High-Level Strategy Pretraining
where $\mathrm{win\_or\_lose}_{j}=1$ if team $j$ wins, $-1$ if it loses, and 00 otherwise; $\mathrm{racket\_hit\_ball}_{j}=1$ if any racket of team $j$ contacts the ball at timestep $t$, and 00 otherwise.
In this stage, we freeze the low-level skills and pretrain the high-level strategy through the team-level self-play method [18] (Fig. 2(a)). We employ an iterative self-play procedure combining population-based evaluation with Nash equilibrium (NE)-based opponent sampling. As the game is symmetric, both teams share a high-level strategy population, and we train only one side against the other. At each iteration, a candidate high-level strategy competes against frozen opponents sampled from the high-level strategy population. Opponent selection follows the NE-based distribution computed from the prior win-rate matrix. When the candidate’s win rate exceeds a threshold or a maximum step count is reached, it is added to the shared population. We then perform pairwise evaluations of all population policies to produce a new win-rate matrix, update the NE-based sampling distribution accordingly, and repeat the cycle. Since the high-level strategy is event-driven, its action steps are temporally sparse. To address this, we apply sample reallocation (Fig. 2(b)), which extracts transitions at event-triggered timesteps and attributes cumulative rewards over the intervals between events. The effectiveness of sample reallocation is analyzed in detail in Sec. 4.3.2.
@@ 4.1 Game Results Against Baselines
Figure 4: Experiment results. (a) HCSP performance against baseline policies. (b) Cross-play win rate heatmap among ten randomly picked policies. (c) Ablation study on policy chaining in Stage I: policy chaining enables successful transition from Attack to Hover, whereas single-policy fails.
@@ 4.2 Analysis of Stage III Co-Self-Play
In this section, we evaluate the impact of Stage III co-self-play within the full training pipeline by comparing the final HCSP policy (i.e., the Stage III policy) with its counterpart trained using only the first two stages (i.e., the Stage II policy), through both quantitative and qualitative analyses.
@@ 4.3.1 Policy Chaining in Stage I
In Stage I, we employ policy chaining, an iterative training method designed to ensure smooth transitions between temporally adjacent low-level skills. We compare policy chaining with a single-policy approach by training the “Attacking the ball to the opponent’s side, followed by Hovering to maintain stability” behavior under the same reward settings. Through policy chaining, we first train an Attack policy using Attack-specific rewards, then train a subsequent Hover policy with Hover-specific rewards. In contrast, the single-policy approach trains one unified policy using the combined Attack and Hover rewards. As shown in Fig. 4(c), both approaches achieve similar performance in the Attack phase. However, only policy chaining successfully learns to transition to Hover after Attack. This demonstrates that policy chaining reduces exploration difficulty by decomposing the complex behavior into sequential, specialized policies.
@@ 4.3.2 High-level Strategy Sample Reallocation in Stage II
Figure 6: Policy population evolution win-rate heatmap in Stage II.
@@ 6 Conclusion
Our approach is currently confined to high-fidelity simulation, which cannot fully reproduce real-world sensor noise, or hardware latency. Consequently, HCSP-trained policies may face a sim-to-real gap and are unlikely to transfer zero-shot to physical drones. Moreover, HCSP assumes access to full state information (drone poses, velocities, and ball dynamics) rather than raw sensor or vision inputs, limiting its applicability in environments where only onboard cameras are available.
@@ Acknowledgments
We sincerely thank Jiayu Chen and Yinuo Chen for their insightful discussions and experimental assistance throughout the course of this work. Their support and thoughtful input have contributed to clarifying our ideas and enhancing the overall quality of the paper.
This research was supported by National Natural Science Foundation of China (No.62406159, 62325405), Postdoctoral Fellowship Program of CPSF under Grant Number (GZC20240830, 2024M761676), China Postdoctoral Science Special Foundation 2024T170496.
@@ A.1 Court
The 3v3 multi‑drone volleyball task is illustrated in Fig. 7(a), with a top‑down, 2D court layout shown in Fig. 7(b). The origin lies at the center of the court, and the y‑axis divides it into two equal halves (each team’s side). The x‑axis runs the 12m length (from $x=-6$ to $x=6$), while the court spans 6m in width along the y‑axis (from $y=-3$ to $y=3$). A net of height 2.43m stretches horizontally along the line $x=0$, from $(0,-3)$ to $(0,3)$.
@@ A.2 Drone
where $\bm{x}_{W}$ and $\bm{v}_{W}$ are the position and velocity of the drone in the world frame, $\bm{R}_{WB}$ is the rotation matrix from the body frame to the world frame, $\bm{J}$ is the diagonal inertia matrix, $\bm{g}$ represents gravity, $\bm{q}$ is attitude in quaternion, and $\bm{\omega}$ is the angular velocity. The operator for quaternion multiplication is denoted by $\otimes$. External forces $\bm{F}$ includes aerodynamic drag and downwash effects. The collective thrust $\bm{f}$ and torque $\bm{\eta}$ are calculated based on the thrust of each rotor $\bm{f}_{i}$ as:
@@ B.1.1 Serve
Observation and reward. The drone’s observation is a vector of dimension $37$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, a one‑hot flag indicating whether the drone is permitted to hit the ball, the ball’s relative position to the target ball position. The detailed specification of the task’s reward function per time step can be found in Table 3.
@@
Before Hit dist_to_ball $[0,0.16]$ drone’s distance to the ball
@@
penalty_pos_z $[-\infty,0]$ being too high or too low
penalty_roll $[-\infty,0]$ excessive roll angle
penalty_yaw $[-\infty,0]$ excessive yaw angle
After Hit dist_to_anchor $[0,16]$ ball’s distance to anchor
End Reward in_side $\{0,10\}$ ball landing in opponent’s court
@@ B.1.2 Receive
Observation and reward. The drone’s observation is a vector of dimension $34$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, a one‑hot flag indicating whether the drone is permitted to hit the ball. The detailed specification of the task’s reward function per time step can be found in Table 4.
@@
Before Hit dist_to_ball $[0,10]$ drone’s distance to the ball
@@
penalty_pos_z $[-\infty,0]$ being too high or too low
penalty_roll $[-\infty,0]$ excessive roll angle
penalty_yaw $[-\infty,0]$ excessive yaw angle
After Hit dist_to_anchor $[0,50]$ ball’s distance to anchor
@@
End Reward in_side $\{0,10\}$ ball landing in opponent’s court
@@ B.1.3 Set
Observation and reward. The drone’s observation is a vector of dimension $34$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, a one‑hot flag indicating whether the drone is permitted to hit the ball. The detailed specification of the task’s reward function per time step can be found in Table 5.
@@
Before Hit dist_to_ball $[0,2]$ drone’s distance to the ball
@@
penalty_pos_xy $[-\infty,0]$ being too far from the ideal point
penalty_pos_z $[-\infty,0]$ being too high or too low
penalty_yaw $[-\infty,0]$ excessive yaw angle
After Hit dist_to_anchor $[0,10]$ ball’s distance to anchor
@@
End Reward highest_ball_pos $[0,400]$ ball’s peak height exceeding 3m
@@ B.1.4 Attack
Observation and reward. The drone’s observation is a vector of dimension $36$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, a one‑hot flag indicating whether the drone is permitted to hit the ball, a one‑hot flag indicating the attack direction. The detailed specification of the task’s reward function per time step can be found in Table 6.
@@
Before Hit dist_to_ball $[0,1]$ drone’s distance to the ball
@@
penalty_yaw $[-2\pi,0]$ excessive yaw angle
penalty_roll $[-2\pi,0]$ excessive roll angle
After Hit dist_to_anchor $[0,150]$ ball’s distance to anchor
@@
End Reward ball_hit_ground $\{0,5\}$ ball hit the ground successfully
ball_final_vel $[0,\infty]$ ball’s velocity hitting the ground
@@ B.1.5 Pass
Observation and reward. The drone’s observation is a vector of dimension $36$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, a one‑hot flag indicating whether the drone is permitted to hit the ball, a one‑hot flag indicating the opponent’s attacking direction. The detailed specification of the task’s reward function per time step can be found in Table 7.
@@
Before Hit dist_to_ball $[0,10]$ drone’s distance to the ball
@@
penalty_pos_z $[-\infty,0]$ being too high or too low
penalty_roll $[-0.5\pi,0]$ excessive roll angle
penalty_yaw $[-0.5\pi,0]$ excessive yaw angle
After Hit dist_to_anchor $[0,10]$ ball’s distance to anchor
@@
End Reward highest_ball_pos $[0,200]$ ball’s peak height exceeding 3m
@@ B.1.6 Hover
Observation and reward. The drone’s observation is a vector of dimension $26$ including the drone’s relative position to the target point, the drone’s root state except position, the drone’s relative orientation to the ideal state. The detailed specification of the task’s reward function per time step can be found in Table 8.
@@
Hover Reward drone_pos $[0,3]$ drone’s relative position to the target point
@@ B.1.7 GoTo
Observation and reward. The drone’s observation is a vector of dimension $26$ including the drone’s relative position to the target point, the drone’s root state except position, the drone’s relative orientation to the ideal state. The detailed specification of the task’s reward function per time step can be found in Table 9.
@@
Hover Reward drone_pos $[0,1]$ drone’s relative position to the target point
@@ B.4 Hyperparameters
The PPO [17] hyperparameters applied to all low‑level skill training tasks are listed in Table 10. All parameters in the table are used unchanged except for the actor and critic learning rates in four cases:
@@
buffer length $64$ num minibatches $16$ PPO epochs $4$
@@
Table 10: Hyperparameters used for PPO in low-level skill training.

### v4

@@ Abstract
In this paper, we tackle the problem of learning to play 3v3 multi-drone volleyball, a new embodied competitive task that requires both high-level strategic coordination and low-level agile control. The task is turn-based, multi-agent, and physically grounded, posing significant challenges due to its long-horizon dependencies, tight inter-agent coupling, and the underactuated dynamics of quadrotors. To address this, we propose Hierarchical Co-Self-Play (HCSP), a hierarchical reinforcement learning framework that separates centralized high-level strategic decision-making from decentralized low-level motion control. We design a three-stage population-based training pipeline to enable both strategy and skill to emerge from scratch without expert demonstrations: (I) training diverse low-level skills, (II) learning high-level strategy via self-play with fixed low-level skills, and (III) joint fine-tuning through co-self-play. Experiments show that HCSP achieves superior performance, outperforming non-hierarchical self-play and rule-based hierarchical baselines with an average 82.9% win rate and a 71.5% win rate against the two-stage variant. Moreover, co-self-play leads to emergent team behaviors such as role switching and coordinated formations, demonstrating the effectiveness of our hierarchical design and training scheme. The project page is at https://sites.google.com/view/hi-co-self-play.
@@
Keywords: Hierarchical Reinforcement Learning, Self-Play, Multi-Agent System
@@ 1 Introduction
To address these challenges, we propose Hierarchical Co-Self-Play (HCSP), a hierarchical reinforcement learning (HRL) framework that separates strategic reasoning from motor control. The policy is decomposed into a high-level strategy responsible for team tactics and multiple low-level skills responsible for drone-specific motion control. The high-level strategy is a centralized multi-layer perceptron (MLP) with three output heads, each producing commands for one drone, enabling synchronized decision-making across the team. The low-level policies are independent across drones and operate at a higher frequency to ensure responsive, fine-grained control. To reflect the turn-based structure of the game, the high-level strategy is only activated upon discrete game events–when the ball is struck or crosses the net–while low-level control runs continuously between these moments.
@@ 3.1 Stage I: Low-Level Skill Acquisition
Pass Pass the ball to the setter Pass from court side $\{\text{left},\ \text{right}\}$
@@
Hover Maintain at a fixed position and altitude Preceding skill $\{\textit{Serve},\textit{Receive},\textit{Pass},\textit{Set},\textit{Attack}\}$
@@ 3.2 Stage II: High-Level Strategy Pretraining
where $\mathrm{win\_or\_lose}_{j}=1$ if team $j$ wins, 1-1 if it loses, and 0 otherwise; $\mathrm{racket\_hit\_ball}_{j}=1$ if any racket of team $j$ contacts the ball at timestep $t$, and 0 otherwise.
In this stage, we freeze the low-level skills and pretrain the high-level strategy through team-level iterative self-play [18] (Fig. 2(a)). As the game is symmetric, both teams share the same high-level strategy population, and we train only one side against the other. At each iteration, a candidate high-level strategy competes against frozen opponents sampled from the population according to the Nash equilibrium (NE)-based distribution derived from the current win-rate matrix. When the candidate’s win rate exceeds a threshold or a maximum step count is reached, it is added to the shared population. We then perform pairwise evaluations of all population policies to produce an updated win-rate matrix, recompute the NE-based sampling distribution, and repeat the cycle. Since the high-level strategy is event-driven, its action steps are temporally sparse, leading to high-variance gradients and unstable learning. To mitigate this issue, we apply sample reallocation (Fig. 2(b)), which extracts transitions at event-triggered timesteps and attributes cumulative rewards over the intervals between events. These event-triggered transitions are stored in a separate buffer for training. The effectiveness of sample reallocation is analyzed in detail in Sec. 4.3.2.
@@ 4.1 Game Results Against Baselines
Figure 4: Experiment results. (a) HCSP performance against baseline policies. (b) Cross-play win rate heatmap among ten randomly picked policies. (c) Ablation study on policy chaining in Stage I: policy chaining enables a successful transition from Attack to Hover, whereas single-policy fails.
@@ 4.3.1 Policy Chaining in Stage I
In Stage I, we employ policy chaining, an iterative training method designed to ensure smooth transitions between temporally adjacent low-level skills. We compare policy chaining with a single-policy approach by training the “Attacking the ball to the opponent’s side, followed by Hovering to maintain stability” behavior under the same reward settings. Through policy chaining, we first train an Attack policy using Attack-specific rewards, then train a subsequent Hover policy with Hover-specific rewards. In contrast, the single-policy approach trains one unified policy using the combined Attack and Hover rewards. In Fig. 4(c), they achieve similar performance in Attack. However, only policy chaining successfully learns Hover after Attack, demonstrating that policy chaining reduces exploration difficulty by decomposing the complex behavior into sequential, specialized policies.
@@ 4.3.2 High-level Strategy Sample Reallocation in Stage II
Figure 6: Comparison of policy population evolution win-rate heatmaps in Stage II.
@@ 6 Conclusion
Our approach is currently limited to high-fidelity simulation, which cannot fully capture real-world sensor noise or hardware latency. Consequently, HCSP-trained policies may encounter a sim-to-real gap and are unlikely to transfer zero-shot to physical drones. Nonetheless, we report preliminary sim-to-real results in Appendix G, which provide promising initial validation in real-world experiments. Furthermore, HCSP assumes access to full state information (drone poses, velocities, and ball dynamics) rather than raw sensor or vision inputs, restricting its applicability in settings where only onboard cameras are available.
@@ Acknowledgments
We sincerely thank Jiayu Chen, Chuqi Wang, and Yinuo Chen for their insightful discussions and experimental assistance throughout the course of this work. We are also grateful to Sicheng He for his initial involvement in the project. Their support and thoughtful input have contributed to clarifying our ideas and enhancing the overall quality of the paper.
This research was supported by National Natural Science Foundation of China (No.62406159, 62325405), Postdoctoral Fellowship Program of CPSF under Grant Number (GZC20240830, 2024M761676), China Postdoctoral Science Special Foundation 2024T170496. Additional support was provided by Tsinghua-Efort Joint Research Center for EAI Computation and Perception, Beijing National Research Center for Information Science and Technology (BNRist), Beijing Innovation Center for Future Chips, and State Key laboratory of Space Network and Communications.
@@ A.1 Court
The 3v3 multi‑drone volleyball task is illustrated in Fig. 7(a), with a top‑down, 2D court layout shown in Fig. 7(b). The origin lies at the center of the court, and the y‑axis divides it into two equal halves (each team’s side). The x‑axis runs the 12m length (from x=6x=-6 to $x=6$), while the court spans 6m in width along the y‑axis (from y=3y=-3 to $y=3$). A net of height 2.43m stretches horizontally along the line $x=0$, from (0,3)(0,-3) to $(0,3)$.
@@ A.2 Drone
where $\bm{x}_{W}$ and $\bm{v}_{W}$ are the position and velocity of the drone in the world frame, $\bm{R}_{WB}$ is the rotation matrix from the body frame to the world frame, $\bm{J}$ is the diagonal inertia matrix, $\bm{g}$ represents gravity, $\bm{q}$ is attitude in quaternion, and $\bm{\omega}$ is the angular velocity. The operator for quaternion multiplication is denoted by $\otimes$. External forces $\bm{F}$ include aerodynamic drag and downwash effects. The collective thrust $\bm{f}$ and torque $\bm{\eta}$ are calculated based on the thrust of each rotor $\bm{f}_{i}$ as:
@@ B.1.1 Serve
Observation and reward. The drone’s observation is a vector of dimension $37$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, a one‑hot flag indicating whether the drone is permitted to hit the ball, and the ball’s relative position to the target ball position. The detailed specification of the task’s reward function per time step can be found in Table 3.
Table 3: Reward of the serve skill
@@
Before
Hit
dist_to_ball $[0,0.16]$ drone’s distance to the ball
@@
penalty_pos_z $\infty,0]$ being too high or too low
penalty_roll $\infty,0]$ excessive roll angle
penalty_yaw $\infty,0]$ excessive yaw angle
After Hit
dist_to_anchor $[0,16]$ ball’s distance to anchor
End
Reward
in_side $\{0,10\}$ ball landing in opponent’s court
@@ B.1.2 Receive
Observation and reward. The drone’s observation is a vector of dimension $34$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, and a one‑hot flag indicating whether the drone is permitted to hit the ball. The detailed specification of the task’s reward function per time step can be found in Table 4.
Table 4: Reward of the receive skill
@@
Before
Hit
dist_to_ball $[0,10]$ drone’s distance to the ball
@@
penalty_pos_z $\infty,0]$ being too high or too low
penalty_roll $\infty,0]$ excessive roll angle
penalty_yaw $\infty,0]$ excessive yaw angle
After
Hit
dist_to_anchor $[0,50]$ ball’s distance to anchor
@@
End
Reward
in_side $\{0,10\}$ ball landing in opponent’s court
@@ B.1.3 Set
Observation and reward. The drone’s observation is a vector of dimension $34$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, and a one‑hot flag indicating whether the drone is permitted to hit the ball. The detailed specification of the task’s reward function per time step can be found in Table 5.
Table 5: Reward of the set skill
@@
Before
Hit
dist_to_ball $[0,2]$ drone’s distance to the ball
@@
penalty_pos_xy $\infty,0]$ being too far from the ideal point
penalty_pos_z $\infty,0]$ being too high or too low
penalty_yaw $\infty,0]$ excessive yaw angle
After
Hit
dist_to_anchor $[0,10]$ ball’s distance to anchor
@@
End
Reward
highest_ball_pos $[0,400]$ ball’s peak height exceeding 3m
@@ B.1.4 Attack
Observation and reward. The drone’s observation is a vector of dimension $36$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, a one‑hot flag indicating whether the drone is permitted to hit the ball, and a one‑hot flag indicating the attack direction. The detailed specification of the task’s reward function per time step can be found in Table 6.
Table 6: Reward of the attack skill
@@
Before
Hit
dist_to_ball $[0,1]$ drone’s distance to the ball
@@
penalty_yaw $2\pi,0]$ excessive yaw angle
penalty_roll $2\pi,0]$ excessive roll angle
After
Hit
dist_to_anchor $[0,150]$ ball’s distance to anchor
@@
End
Reward
ball_hit_ground $\{0,5\}$ ball hit the ground successfully
ball_final_vel $0,\infty]$ ball’s velocity hitting the ground
@@ B.1.5 Pass
Observation and reward. The drone’s observation is a vector of dimension $36$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, a one‑hot flag indicating whether the drone is permitted to hit the ball, and a one‑hot flag indicating the opponent’s attacking direction. The detailed specification of the task’s reward function per time step can be found in Table 7.
Table 7: Reward of the pass skill
@@
Before
Hit
dist_to_ball $[0,10]$ drone’s distance to the ball
@@
penalty_pos_z $\infty,0]$ being too high or too low
penalty_roll $0.5\pi,0]$ excessive roll angle
penalty_yaw $0.5\pi,0]$ excessive yaw angle
After
Hit
dist_to_anchor $[0,10]$ ball’s distance to anchor
@@
End
Reward
highest_ball_pos $[0,200]$ ball’s peak height exceeding 3m
@@ B.1.6 Hover
Observation and reward. The drone’s observation is a vector of dimension $26$ including the drone’s relative position to the target point, the drone’s root state except position, and the drone’s relative orientation to the ideal state. The detailed specification of the task’s reward function per time step can be found in Table 8.
Table 8: Reward of the hover skill
@@
Hover
Reward
drone_pos $[0,3]$ drone’s relative position to the target point
@@ B.1.7 GoTo
Observation and reward. The drone’s observation is a vector of dimension $26$ including the drone’s relative position to the target point, the drone’s root state except position, and the drone’s relative orientation to the ideal state. The detailed specification of the task’s reward function per time step can be found in Table 9.
Table 9: Reward of the goto skill
@@
Hover
Reward
drone_pos $[0,1]$ drone’s relative position to the target point
@@ B.4 Hyperparameters
The PPO [17] hyperparameters applied to all low‑level skill training tasks are listed in Table 10. All parameters in the table are used unchanged except for the actor and critic learning rates:
@@
Table 10: Hyperparameters used for PPO in low-level skill training.
@@
buffer length $64$ num minibatches $16$ PPO epochs 44